Let X1, X2, ...,Xn X , X be a random sample from a distribution with density function
f(x,θ) = {θ^-1 e^-x/θ, if X>0
{0 , otherwise
Show that X=n^-1 ∑Xi is unbiased for θ .
1
Expert's answer
2019-04-04T10:01:35-0400

The first moment of X is 


"\mu_1=E(X)=\displaystyle\intop_{0}^\infin{x \over \theta}e^{-x / \theta}dx"

"\int udv=uv-\int vdu"

"u={x\over\theta}, du={1\over\theta}dx, v=-\theta e^{-x/ \theta}"

"\int {x \over \theta}e^{-x / \theta}dx=-xe^{-x/\theta}+\int e^{-x/\theta}dx=-xe^{-x/\theta}-\theta e^{-x/\theta}+C"

"\mu_1=E(X)=\displaystyle\intop_{0}^\infin{x \over \theta}e^{-x / \theta}dx=[-xe^{-x/\theta}-\theta e^{-x/\theta}]\begin{matrix}
 \infin \\
 0
\end{matrix}=\theta"

Equating the sample first moment to the population first moment:  


"\mu_1=\widehat{\mu}_1"

"\theta={1\over n}\displaystyle\sum_{i=1}^nX_i=\bar{X}"

Therefore,


"\widehat{\theta}={1\over n}\displaystyle\sum_{i=1}^nX_i=\bar{X}"

"{1\over n}\displaystyle\sum_{i=1}^nX_i=\bar{X}"  2 is an unbiased estimator of "\theta".